# NVS-SQA: Exploring Self-Supervised Quality Representation Learning for Neurally Synthesized Scenes without References

### v1

@@ Abstract
Index Terms:
Perceptual Quality Assessment, Quality of Experience (QoE), Immersive Experience, No-Reference Quality Assessment, Self-Supervised Learning, Novel View Synthesis, 3D Reconstruction, 3D Gaussian Splatting, Neural Radiance Fields (NeRF).
@@ I Introduction
To address these challenges, we introduce NVS-SQA, the first self-supervised learning framework for NSS quality assessment, offering distinct advantages as illustrated in Fig. 3. This framework aims to derive effective no-reference quality representations for NSS from unlabeled datasets. NVS-SQA leverages heuristic cues and quality scores as soft learning objectives to calibrate the distance between the quality representations of contrastive pair (detailed in Section III-B). To converges toward an optimal multi-branch guidance for quality representation learning, we design an adaptive weighting method inspired from the statistical principles. Furthermore, we develop a deep neural network as the backbone of our framework, designed to handle scenes of varying view counts and spatial resolutions. In the experiments, the proposed method was trained on an unlabeled dataset. Subsequently, its weights were frozen, and it was evaluated using linear regression on three labeled datasets comprising various new scenes produced by entirely unseen NVS methods (i.e., entirely distinct from the training dataset). The results demonstrate that the proposed approach surpasses 17 existing no-reference quality assessment methods in assessing NSS and even outperforms 16 mainstream full-reference quality assessments across six out of six evaluation metrics. The key points of this paper are summarized as follows:
@@ II Related work
Figure 4: Overview of the proposed learning framework. The framework consists of two main stages: (1) The self-supervised quality representation learning stage, where AdaptiSceneNet (described in Section III-D) learns quality representations from unlabeled NSS through a contrastive pair preparation (detailed in Section III-C) process and multi-branch guidance adaptation for quality representation learning (described in Section III-B), and (2) The perceptual quality estimation stage, where the pretrained AdaptiSceneNet is used to estimate perceptual quality by mapping learned representations to human scores.
@@ III Methodology
III-A Method overview
@@
The self-supervised quality representation learning stage aims to train a neural network capable of generating effective quality representations without the need for reference views or human perceptual annotations. During this stage, unlabeled NSS undergo a contrastive pair preparation process (elaborated in Section III-C) to create diverse pairs for subsequent training. These pairs are then feed to AdaptiSceneNet, a neural network designed for compatibility with varying scenes (detailed in Section III-D), which processes the pairs separately but under shared weights to produce quality representation pairs. We introduce a multi-branch guidance adaptation (described in Section III-B) to direct the quality representation learning process.
@@
III-B Multi-branch guidance adaptation for quality representation learning
@@
To address these issues, we design multiple objectives for quality representation learning. The key idea is to utilize quality scores from various full-reference quality assessment methods, along with the proportion of replaced views in an NSS as complementary heuristic cues (detailed in Section III-C). Different from widely used discrete positive/negative labels for contrastive pairs [47, 49, 29, 30], this approach provides continuous and comprehensive targets, enhancing the effectiveness of learning. Another advantage of this approach is that, unlike popular InfoNCE-like loss functions [48, 47, 49, 29, 30], it does not require loading a batch of negative pairs alongside a positive pair for each gradient update step (as shown in the equation 1), and hence reduce memory usage. This efficiency is particularly relevant in the context of the extensive data volumes associated with NSS.
A workflow of the proposed multi-branch guidance adaptation is illustrated in the right segment of Fig. 4. The quality representations are passed through three distinct non-linear projectors: an IQA-guided branch, a VQA-guided branch, and a replacement ratio branch. The IQA-guided branch utilizes insights from full-reference IQA methods to assess the static quality of NSS, while the VQA-guided branch employs full-reference VQA methods for evaluating dynamic quality aspects. Specifically, a lower IQA or VQA score for a contrastive pair implies a greater disparity in quality. The replacement ratio, termed as REP, serves as a complementary heuristic, suggesting that a higher proportion of view replacements in NSS likely results in more distinct representations. Consequently, our objective function is formulated to reflect that greater differences in replacement ratios, VQA scores, or IQA scores should correspond to more divergent representations (i.e., less similarity). Formally, given a set of contrastive pairs, $\{(s_{1}^{i},s_{2}^{i})\}_{i=1}^{N}$, and denoting the model as $H$, we define each branch, $\psi\in\Psi:=\{IQA,\;VQA,\;REP\}$, of learning loss as follows:
$\ell_{\psi}^{i}(H)=\left|\;sim\left[\tau_{\psi}\cdot H(s_{1}^{i}),\;\tau_{\psi}\cdot H(s_{2}^{i})\right]-\;\psi(s_{1}^{i},s_{2}^{i})\;\right|,$ (1)
where $sim(u,v)=u^{T}v/\|u\|\|v\|$ denotes the cosine similarity between $l_{2}$ normalized vectors $u$ and $v$, and $\tau_{\psi}$ represents the non-linear MLP projector [47] corresponding to each branch. The output from IQA, VQA methods, and the replacement ratio, REP, are rescaled to the cosine similarity value range of $[-1,1]$.
@@
where $\lambda_{\psi}$ denotes the weight assigned to each loss component, adjustable through tuning. This formulation underscores our strategy to adaptively learn quality representations from NSS, leveraging a combination of heuristic cues and quality assessment insights. We performed a grid search over a predefined search space to identify the optimal weights, as demonstrated in the ablation studies in Section IV-D. The optimal results of MBW are also presented in the experimental results in Section IV.
@@
$p(\psi(s_{1},s_{2})|\delta_{H}(s_{1},s_{2}))=\mathcal{N}\left(\delta_{H}(s_{1},s_{2}),\sigma_{\psi}^{2}\right),$ (3)
where $\delta_{H}(s_{1},s_{2}):=sim\left[\tau_{\psi}\cdot H(s_{1}),\;\tau_{\psi}\cdot H(s_{2})\right]$ represents the estimated similarities between the embeddings of contrastive pairs $\{(s_{1},s_{2})\}$ given a representation model $H$, and $\sigma_{\psi}$ is the observed branch-dependent noise. The multi-branch probabilistic model is then given by:
$p(\Psi|\delta_{H})$ $=p(\{\psi(s_{1},s_{2})\}^{\Psi}|\delta_{H}(s_{1},s_{2}))$ (4)
$=\prod_{\psi}^{\Psi}{p(\psi(s_{1},s_{2})|\delta_{H}(s_{1},s_{2}))}$
$=\prod_{\psi}^{\Psi}{\mathcal{N}\left(\delta_{H}(s_{1},s_{2}),\sigma_{\psi}^{2}\right)}.$
@@
$\log p(\Psi|\delta_{H})\propto-\sum_{\psi}^{\Psi}\biggl{(}\frac{1}{2\sigma_{\psi}^{2}}\left||\psi(s_{1},s_{2})-\delta_{H}(s_{1},s_{2})\right||^{2}$ (5)
$+\log\sigma_{\psi}\biggl{)}.$
The learning objective can then be formulated as minimizing the negative log-likelihood:
$H^{*}=\arg\min_{H,\sigma}\frac{1}{N}\sum_{i=1}^{N}\sum_{\psi\in\Psi}\tilde{\ell}_{\psi}^{i}(H,\sigma_{\psi}),$ (6)
where the loss function for each branch is define as:
~$\tilde{\ell}_{\psi}^{i}(H,\sigma_{\psi})=\frac{1}{2\sigma_{\psi}^{2}}\biggl{|}\;\text{sim}\left(\tau_{\psi}\cdot H(s_{1}^{i}),\;\tau_{\psi}\cdot H(s_{2}^{i})\right)$ (7)
$-\psi(s_{1}^{i},s_{2}^{i})\;\biggr{|}^{2}+\log\sigma_{\psi},$
where $\sigma_{\Psi}$ represents the learnable noise parameter for each branch, the term $\frac{1}{2\sigma_{\Psi}^{2}}$ dynamically scales the loss of each branch based on its noise, and the term $\log\sigma_{\Psi}$ serves as a regularizer to prevent $\sigma_{\Psi}$ from becoming too large.
@@
III-C Contrastive pair preparation
@@
The preparation of pairs is crucial for acquiring robust representations, particularly in semantic representation learning. For instance, SimCLR [47] utilizes data augmentation techniques such as random cropping, color distortion, and Gaussian blur, designating crops from the same image instance as positive pairs and others as negative. However, these methods may not be sufficient within the context of NSS. Given the limited availability of NSS, the likelihood of generalizing self-supervised learned representations is reduced, highlighting the need for a more varied set of contrastive pairs from a restricted dataset. As discussed in Section III-B and illustrated in Fig. 5, the “same instance, similar representation” assumption proves unreliable for quality representation learning in NSS, making traditional positive/negative partitioning inappropriate.
@@
To address these challenges, we have developed an algorithm for the preparation of contrastive pairs specifically tailored to NSS, as depicted in the leftmost block of Fig. 4. According to the objectives outlined in Section III-B, we do not classify pairs as positive or negative, allowing us to create NSS pairs at any quality distance and maximize diversity by applying varying levels of distortion. We further randomize crops and clips to ensure the model accommodates NSS with different numbers of views and spatial resolutions. Additionally, we randomly replace views with those from another NVS method to enhance variability. Importantly, to facilitate the effectiveness of quality representation learning, we ensure that each NSS pair represents the same scene, thereby controlling for semantic consistency.
@@
Formally, the proposed pair preparation algorithm is outlined in Algorithm 1 titled Preparing Contrastive Pairs. Each pair, maintaining identical semantic information, captures variations in scene quality through controlled distortions and modifications. Starting with an unlabeled dataset $D$ and a target of $N$ pairs, the algorithm creates a set $P$ of pairs. It selects an NSS randomly from $D$, applies random cropping and orientation adjustments to form a base scene $s_{1}$, and then duplicates, distorts, and partially replaces views in $s_{1}$ with those generated by an alternate NVS method to produce its counterpart $s_{2}$. These steps introduce necessary variability and contrast, essential for training models to discern subtle visual differences. The algorithm also adjusts the number of views and spatial resolutions randomly, ensuring the model’s compatibility with various NSS sizes. The replacement ratio, REP, moderates the disparity between $s_{1}$ and $s_{2}$, serving as a heuristic cue for self-supervised learning (as detailed in Section III-B). By repeatedly forming such pairs, the algorithm compiles a diverse set $P$ ready for subsequent self-supervised learning.
@@
1:Unlabeled dataset $D$ of NSS, number of result pairs $N$.
@@
11:    $s_{2}\leftarrow\text{ReplaceViews}(s_{2},\text{AnotherRandomNVS}(views),\text{REP})$

### v3

@@ I Introduction
To address these challenges, we introduce NVS-SQA, the first self-supervised learning framework for NSS quality assessment, offering distinct advantages as illustrated in Fig. 3. This framework aims to derive effective no-reference quality representations for NSS from unlabeled datasets. NVS-SQA leverages heuristic cues and quality scores as soft learning objectives to calibrate the distance between the quality representations of contrastive pair (detailed in Section III-C). To converges toward an optimal multi-branch guidance for quality representation learning, we design an adaptive weighting method inspired from the statistical principles. Furthermore, we develop a deep neural network as the backbone of our framework, designed to handle scenes of varying view counts and spatial resolutions. In the experiments, the proposed method was trained on an unlabeled dataset. Subsequently, its weights were frozen, and it was evaluated using linear regression on three labeled datasets comprising various new scenes produced by entirely unseen NVS methods (i.e., entirely distinct from the training dataset). The results demonstrate that the proposed approach surpasses 17 existing no-reference quality assessment methods in assessing NSS and even outperforms 16 mainstream full-reference quality assessments across six out of six evaluation metrics. The key points of this paper are summarized as follows:
@@ II Related work
Figure 4: Overview of the proposed learning framework. The framework consists of two main stages: (1) The self-supervised quality representation learning stage, where AdaptiSceneNet (described in Section III-E) learns quality representations from unlabeled NSS through a contrastive pair preparation (detailed in Section III-D) process and multi-branch guidance adaptation for quality representation learning (described in Section III-C), and (2) The perceptual quality estimation stage, where the pretrained AdaptiSceneNet is used to estimate perceptual quality by mapping learned representations to human scores.
@@ III Methodology
III-A Problem Definition
Let $S=\{s^{1},s^{2},...,s^{n}\}$ be a neurally synthesized scene (NSS) composed of $n$ views generated by a Novel View Synthesis (NVS) method. The goal of Neurally Synthesized Scene Quality Assessment (NSS-QA) is to learn a function $f:S\rightarrow q\in\mathbb{R}$ where $q$ is a scalar score that reflects the perceptual quality of the scene as perceived by human observers. This score should account for comprehensive quality of NSS, including spatial fidelity, angular and temporal coherence, and structure integrity, capturing distortions that occur within individual views as well as inconsistencies across viewpoints. In the no-reference setting, the function $f$ must estimate perceptual quality without access to any ground-truth or reference views, relying solely on the synthesized scene itself. This makes the problem particularly challenging, as the model must infer perceptual degradation patterns directly from generated content. The predicted quality score is expected to satisfy three key properties: (1) strong correlation with human subjective ratings across varied scenes and synthesis methods, (2) sensitivity to perceptual artifacts including spatial degradations, temporal or angular inconsistencies, and geometric misalignments, and (3) generalization across scenes with diverse resolutions, view counts, and rendering pipelines. In this work, we address this problem through a self-supervised learning approach that leverages perceptually motivated guidance signals and contrastive representation learning to develop a robust no-reference quality estimator.
III-B Method overview
@@
The self-supervised quality representation learning stage aims to train a neural network capable of generating effective quality representations without the need for reference views or human perceptual annotations. During this stage, unlabeled NSS undergo a contrastive pair preparation process (elaborated in Section III-D) to create diverse pairs for subsequent training. These pairs are then feed to AdaptiSceneNet, a neural network designed for compatibility with varying scenes (detailed in Section III-E), which processes the pairs separately but under shared weights to produce quality representation pairs. We introduce a multi-branch guidance adaptation (described in Section III-C) to direct the quality representation learning process.
@@
III-C Multi-branch guidance adaptation for quality representation learning
@@
To address these issues, we design multiple objectives for quality representation learning. The key idea is to utilize quality scores from various full-reference quality assessment methods, along with the proportion of replaced views in an NSS as complementary heuristic cues (detailed in Section III-D). Different from widely used discrete positive/negative labels for contrastive pairs [47, 49, 29, 30], this approach provides continuous and comprehensive targets, enhancing the effectiveness of learning. Another advantage of this approach is that, unlike popular InfoNCE-like loss functions [48, 47, 49, 29, 30], it does not require loading a batch of negative pairs alongside a positive pair for each gradient update step (as shown in the equation 1), and hence reduce memory usage. This efficiency is particularly relevant in the context of the extensive data volumes associated with NSS.
The perceived quality of NSS is primarily influenced by three key factors: spatial fidelity, angular and temporal coherence, and structural integrity. To address these dimensions, we propose a multi-branch guidance adaptation strategy, illustrated in the right segment of Fig. 4. In this design, the learned quality representations are passed through three distinct non-linear projectors, each guided by a different supervisory signal: an IQA-guided branch, a VQA-guided branch, and a replacement ratio (REP) branch. The IQA-guided branch leverages full-reference image quality assessment metrics to model spatial fidelity within individual views. The VQA-guided branch incorporates full-reference video quality assessment metrics to capture angular and temporal coherence across views. The REP branch provides a complementary cue based on the view replacement ratio, which reflects the proportion of synthesized views that are substituted due to occlusions or reconstruction defects. This ratio is typically indicative of structural degradation within the scene. Together, these branches provide complementary quality cues: the IQA branch emphasizes static spatial details, the VQA branch accounts for dynamic artifacts across views, and the REP branch captures structural inconsistencies and geometric distortions unique to multi-view synthesis. This multi-branch design enables the model to learn a robust and perceptually aligned embedding space.
To ensure notation consistency, we summarize the key symbols as follows:
  • $(s_{1}^{i},s_{2}^{i})$: the $i$-th contrastive pair of NSS samples;
  • $\Psi:=\{\text{IQA},\text{VQA},\text{REP}\}$: the set of guidance branches;
  • $H$: the shared representation network;
  • $\tau_{\psi}$: branch-specific MLP projection for $\psi$;
  • $\psi(s_{1}^{i},s_{2}^{i})$: the rescaled guidance signal (in [1,1][-1,1]) for branch $\psi$ on pair $(s_{1}^{i},s_{2}^{i})$;
  • $\text{sim}(u,v)$: cosine similarity, defined as $\frac{u^{\top}v}{\|u\|\|v\|}$;
Formally, given a set of contrastive pairs, $\{(s_{1}^{i},s_{2}^{i})\}_{i=1}^{N}$, and a representation model $H$, we define each guidance branch $\psi\in\Psi$, with branch-specific projector $\tau_{\psi}$, and define the learning loss for each pair as:
$\ell_{\psi}^{i}(H)=\left|\;\text{sim}\left(\tau_{\psi}H(s_{1}^{i}),\;\tau_{\psi}H(s_{2}^{i})\right)\;-\;\psi(s_{1}^{i},s_{2}^{i})\;\right|,$ (1)
where $\text{sim}(u,v)=\frac{u^{\top}v}{\|u\|\|v\|}$ denotes the cosine similarity between $l_{2}$-normalized vectors $u$ and $v$. The outputs of the IQA, VQA, and REP branches are all linearly rescaled to the range [1,1][-1,1] to match the similarity scale. Specifically, for a given scalar quality score $x$ in the range $[a,b]$, the rescaled value $\tilde{x}$ is computed as $\tilde{x}=2\cdot\frac{x-a}{b-a}-1$. The values of $a$ and $b$ are defined empirically based on the score distribution in the training set for each branch.
It is important to note that the three guidance cues serve as soft and complementary supervision signals to guide the learning of quality-aware representations. The purpose is not to replicate the behavior of any individual cue but to leverage their distinct perspectives to shape a representation space that aligns more closely with human perceptual judgments. Each of the three guidance branches corresponds to a distinct aspect of perceptual quality. The IQA-guided branch captures spatial fidelity within individual synthesized views, where artifacts such as blurring or texture inconsistencies commonly occur. The VQA-guided branch emphasizes angular and temporal coherence across viewpoints, which helps detect issues like flickering, reprojection errors, or perceptual instability that arise from view-to-view inconsistencies. The REP-guided branch focuses on structural integrity by reflecting disruptions introduced through controlled view substitutions, such as geometric misalignments or parallax distortions. These three guidance branches act as soft supervision signals rather than strict labels, allowing the model to align the representational distance with perceptual degradations across a range of synthesis artifacts. By jointly leveraging all the three branches, NVS-SQA learns a more comprehensive and perceptually robust quality-aware embedding space, which is further supported by the quantitative and qualitative improvements observed in the ablation study (Section IV-D).
While our approach adopts cosine similarity as a contrastive objective, it diverges significantly from prior contrastive learning methods such as SimCLR [47]. SimCLR is designed for semantic representation learning from individual images, where the assumption is that different augmentations of the same image share the same semantics. In contrast, our approach operates on multi-view NSS, where the goal is not to preserve semantics but to learn perceptual quality embeddings that reflect not only intra-view quality but also view-dependent distortions, such as geometric inconsistencies, occlusions, and synthesis artifacts. Cosine similarity is retained in our formulation not to mimic SimCLR, but because it provides a scale-invariant, bounded similarity metric that remains effective under the variability of view count and spatial resolution in novel view synthesis. In our context, cosine distance serves as a perceptual proxy between multi-view NSS clips, enabling the model to learn quality-aware representations that are sensitive to structural degradations and cross-view inconsistencies.
@@
where $\lambda_{\psi}$ denotes the weight assigned to each branch-specific loss component, which can be adjusted via grid search (see ablation in Section IV-D). This formulation adaptively combines heuristic guidance cues to optimize the quality representation.
@@
$p(\psi(s_{1}^{i},s_{2}^{i})|\delta_{H,\psi}(s_{1}^{i},s_{2}^{i}))=\mathcal{N}\left(\delta_{H,\psi}(s_{1}^{i},s_{2}^{i}),\sigma_{\psi}^{2}\right),$ (3)
where $\delta_{H,\psi}(s_{1}^{i},s_{2}^{i}):=\text{sim}\left(\tau_{\psi}H(s_{1}^{i}),\;\tau_{\psi}H(s_{2}^{i})\right)$ represents the predicted similarity for branch $\psi$, and $\sigma_{\psi}$ is the learnable noise parameter for branch $\psi$.
The joint multi-branch probabilistic model is then:
$p(\{\psi(s_{1}^{i},s_{2}^{i})\}_{\psi\in\Psi}|\{\delta_{H,\psi}(s_{1}^{i},s_{2}^{i})\})=\prod_{\psi\in\Psi}{\mathcal{N}\left(\delta_{H,\psi}(s_{1}^{i},s_{2}^{i}),\sigma_{\psi}^{2}\right)}.$ (4)
@@
$\log p(\{\psi(s_{1}^{i},s_{2}^{i})\}_{\psi\in\Psi}|\{\delta_{H,\psi}(s_{1}^{i},s_{2}^{i})\})\propto$ (5)
$-\sum_{\psi\in\Psi}\left[\frac{1}{2\sigma_{\psi}^{2}}\left(\psi(s_{1}^{i},s_{2}^{i})-\delta_{H,\psi}(s_{1}^{i},s_{2}^{i})\right)^{2}+\log\sigma_{\psi}\right].$
The learning objective is then minimizing the negative log-likelihood over all pairs:
$H^{*}=\arg\min_{H,\{\sigma_{\psi}\}}\frac{1}{N}\sum_{i=1}^{N}\sum_{\psi\in\Psi}\tilde{\ell}_{\psi}^{i}(H,\sigma_{\psi}),$ (6)
where the loss function for each branch is defined as:
$\tilde{\ell}_{\psi}^{i}(H,\sigma_{\psi})=\frac{1}{2\sigma_{\psi}^{2}}\left|\;\text{sim}\left(\tau_{\psi}H(s_{1}^{i}),\;\tau_{\psi}H(s_{2}^{i})\right)-\psi(s_{1}^{i},s_{2}^{i})\right|^{2}$ (7)
$+\log\sigma_{\psi},$
where each $\sigma_{\psi}$ dynamically scales its respective loss branch and regularizes the overall loss, ensuring balanced training across IQA, VQA, and REP.
@@
III-D Contrastive pair preparation
@@
The preparation of pairs is crucial for acquiring robust representations, particularly in semantic representation learning. For instance, SimCLR [47] utilizes data augmentation techniques such as random cropping, color distortion, and Gaussian blur, designating crops from the same image instance as positive pairs and others as negative. However, these methods may not be sufficient within the context of NSS. Given the limited availability of NSS, the likelihood of generalizing self-supervised learned representations is reduced, highlighting the need for a more varied set of contrastive pairs from a restricted dataset. As discussed in Section III-C and illustrated in Fig. 5, the “same instance, similar representation” assumption proves unreliable for quality representation learning in NSS, making traditional positive/negative partitioning inappropriate.
@@
To address these challenges, we have developed an algorithm for the preparation of contrastive pairs specifically tailored to NSS, as depicted in the leftmost block of Fig. 4. According to the objectives outlined in Section III-C, we do not classify pairs as positive or negative, allowing us to create NSS pairs at any quality distance and maximize diversity by applying varying levels of distortion. We further randomize crops and clips to ensure the model accommodates NSS with different numbers of views and spatial resolutions. Additionally, we randomly replace views with those from another NVS method to enhance variability. Importantly, to facilitate the effectiveness of quality representation learning, we ensure that each NSS pair represents the same scene, thereby controlling for semantic consistency.
@@
Formally, the proposed pair preparation algorithm is outlined in Algorithm 1 titled Preparing Contrastive Pairs. Each pair, maintaining identical semantic information, captures variations in scene quality through controlled distortions and modifications. Starting with an unlabeled dataset $D$ and a target of $N$ pairs, the algorithm creates a set $P$ of pairs. It selects an NSS randomly from $D$, applies random cropping and orientation adjustments to form a base scene $s_{1}$, and then duplicates, distorts, and partially replaces views in $s_{1}$ with those generated by an alternate NVS method to produce its counterpart $s_{2}$. These steps introduce necessary variability and contrast, essential for training models to discern subtle visual differences. The algorithm also adjusts the number of views and spatial resolutions randomly, ensuring the model’s compatibility with various NSS sizes. The replacement ratio, $r$, moderates the disparity between $s_{1}$ and $s_{2}$, serving as a heuristic cue for self-supervised learning (as detailed in Section III-C). In our implementation, the replacement ratio $r$ is randomly sampled from a uniform distribution $r\sim\mathcal{U}(0,0.5)$. This ratio controls the percentage of views in a base scene $s_{1}$ that are replaced with views synthesized by a different NVS method to create $s_{2}$. Through empirical exploration, the range $[0,0.5]$ allows for diverse and controllable quality differences while avoiding redundancy, since replacement ratios beyond $0.5$ are equivalent to swapping the roles of $s_{1}$ and $s_{2}$. All replacements are performed within the same scene to maintain semantic consistency, and the replacement methods are selected randomly to ensure balance and reduce bias. By repeatedly forming such pairs, the algorithm compiles a diverse set $P$ ready for subsequent self-supervised learning. Moreover, there are several designs to prevent unintended biases from disparities across NVS methods. For example, all view replacements are restricted to the same scene to ensure content alignment. The substituting NVS methods are selected uniformly at random during training to avoid overexposure to any single method. Additionally, the model is not supervised to identify or rank specific methods. Instead, soft quality cues, including image-level (IQA), video-level (VQA), and structural (REP) signals, are used to guide representation learning. These cues enable the model to learn perceptual quality embeddings that generalize beyond individual method characteristics.
To analyze the scalability of the proposed contrastive pair preparation, we consider a training setup with $s$ scenes, each containing $v$ distinct view sequences (i.e., consistent camera paths or spatial neighborhoods). For each view sequence, $m$ different NVS methods generate corresponding multi-view clips. From each NVS rendering, $c$ clips are extracted through randomized sampling and pre-processing. This results in $mc$ clips per view sequence. Contrastive pairs are only formed between clips corresponding to the same view sequence, ensuring semantic alignment while allowing variation across NVS methods and augmentations. The number of unordered clip pairs per view sequence is $P_{\text{view}}=\binom{mc}{2}=[mc(mc-1)]/2$. The total number of base contrastive pairs across all view sequences and scenes is $P_{\text{total}}=s\cdot v\cdot[mc(mc-1)]/2$. Each contrastive pair can then be further diversified using $a$ augmentation variants, including geometric transformations, appearance distortions, resolution changes, and view-level replacements. The effective number of unique training signals becomes $P_{\text{effective}}=a\cdot s\cdot v\cdot[mc(mc-1)]/2$. For example, using conservative values such as $s=10$ scenes, $v=300$ view sequences per scene, $m=5$ NVS methods, $c=5$ clips per method, and $a=20$ augmentation variants, the total effective training signal reaches tens of millions of unique contrastive pairs. This demonstrates that even under a limited number of scenes, the proposed pair preparation strategy enables large-scale self-supervised learning through combinatorial growth.
@@
1:Unlabeled dataset $D$ of NSS, number of result pairs NN.
@@
11:  $s_{2}\leftarrow\text{ReplaceViews}(s_{2},\text{AnotherRandomNVS}(views),r)$